OpenStudy (anonymous):

2. Yves has bought 420 tropical fish for a museum display. He bought 6 times as many parrotfish as angelfish. How many of each type of fish did he buy? (Points : 1) 30 angelfish, 180 parrotfish 60 angelfish, 360 parrotfish 120 angelfish, 200 parrotfish 360 angelfish, 60 parrotfish

OpenStudy (hockeychick23):

60 angelfish and 360 parrotfish

OpenStudy (hockeychick23):

60+60(6)= 60+360=420